OpenStudy (anonymous):

du/dt = (u^2+4)/(t^2+4) I got to the point where I have: (tan(u/2))^-1=(tan(t/2))^-1 +c but I don't know where to go from there. Please help!

OpenStudy (anonymous):

Take the tangent of both sides: \[\frac{u}{2}=\tan\left(\tan^{-1}\frac{t}{2}+C\right)\] Use the angle sum identity for tangent: \[\tan(x+y)=\frac{\tan x+\tan y}{1-\tan x\tan y}\] This gives \[\frac{u}{2}=\frac{\tan\left(\tan^{-1}\dfrac{t}{2}\right)+\tan C}{1-\tan\left(\tan^{-1}\dfrac{t}{2}\right)\tan C}\] The tangent of a constant is another constant, so you reduce to \[\frac{u}{2}=\frac{\tan\left(\tan^{-1}\dfrac{t}{2}\right)+C}{1-C\tan\left(\tan^{-1}\dfrac{t}{2}\right)}\] Since \(\tan x\) and \(\tan^{-1}x\) are inverse functions, you have \(\tan\left(\tan^{-1}x\right)=x\), which gives you \[\frac{u}{2}=\frac{\dfrac{t}{2}+C}{1-C\dfrac{t}{2}}\] Some simplifying gives \[\frac{u}{2}=\frac{t+2C}{2-Ct}\] Twice a constant is another constant: \[\frac{u}{2}=\frac{t+C}{2-Ct}\] And isolate \(u\): \[u=\frac{2t+C}{2-Ct}\]

OpenStudy (anonymous):

The constant rule will let you simplify further, actually, to \[u=\frac{t+C}{1-Ct}\]